# The Landscape of Fairness: An Axiomatic and Predictive Framework for Network QoE Sensitivity

### v1

@@ Abstract
Index Terms:
QoE, Fairness, Entropy, Logistic Weighting, Network Robustness
@@ I Introduction
This paper develops such a framework, centered on an axiomatically-derived QoE-Imbalance metric ($I$) that captures functional fairness through a tunable Sigmoid function ($s(h;a,h_{0})$). Our contributions are threefold:
@@
Ultimately, this work transforms fairness from a static score into a predictable and navigable landscape, offering a new design philosophy for engineering more fundamentally robust and resilient network systems.
@@ II-A An Axiomatic Foundation for Imbalance
Theorem A.1 (Uniqueness of the Imbalance Metric).
@@ II-C Analytical Properties and Invariance
A prerequisite for our calculus-based sensitivity analysis is that the objective functions are well-behaved. We formally establish that the metrics $I(a,h_{0})$ and $\bar{s}(a,h_{0})$ are both continuous and differentiable with respect to the SLA parameters $a$ and $h_{0}$ across their entire domain (see Appendices B for formal proofs). This ensures our framework is robust and validates the use of gradients and Hessians.
@@
The proof follows directly from the fact that the exponent in the Sigmoid function, $a(h_{i}-h_{0})$, remains unchanged under both transformations. Since all weights $\{w_{u,v}\}$ are invariant, the resulting metrics $I$ and $\bar{s}$ are also invariant. ∎
@@ III-C Second-Order Analysis: The Hessian as a Risk Detector
In the differentiable domain described in Appendix C, the mixed second-order partial derivatives are equal due to Clairaut’s theorem, making the Hessian symmetric. Specifically, the term $|\partial^{2}I/\partial h_{0}^{2}|$ serves as a powerful, quantitative “cliff detector,”…
@@ III-D A Unified Methodology for Sensitivity Analysis
    • Area of Robustness (AoR): The total area of the Optimal Operating Region. A larger AoR signifies greater resilience and adaptability to a wider range of service demands.
    • Maximum Curvature Risk (MCR): The maximum value of the curvature ($|\partial^{2}I/\partial h_{0}^{2}|$) found along the boundary of the Optimal Operating Region. This quantifies the ”sharpness of the cliff” at the edge of acceptable performance, with lower values indicating a safer design.
Mathematically, assuming the optimal operating region $R$ is defined by a set of constraints (e.g., $I\leq I^{*}$ and $\bar{s}\geq\bar{s}^{*}$) in the $(a,h_{0})$ space, these metrics are calculated as:
  • AoR is the area of this region, approximated in practice by summing the areas of the discrete grid cells that satisfy the constraints:
    $\text{AoR}=\iint_{R}da\,dh_{0}$ (11)
  • MCR is the maximum value of the curvature with respect to $h_{0}$ found along the boundary $\partial R$ of this optimal region:
@@ IV-B1 Validation of the Small-a Second-Order Law
Theorem 3 predicts that for tolerant SLAs ($a\to 0$), the imbalance $I$ is quadratically proportional to the strictness $a$ ($I\propto a^{2}$), with a coefficient determined by the network’s path-length variance, $\mathrm{Var}(h)$. Figure 1 provides a direct test of this prediction. As can be seen, the simulation data for all four topologies closely follow a linear trend when $I$ is plotted against $a^{2}$, confirming the quadratic relationship. The slopes of these lines, however, vary dramatically, indicating a strong dependence on the underlying topology. The Path graph, with its extremely dispersed path lengths, is by far the most sensitive, while the ultra-compact Star graph is the least.
@@ IV-B2 Validation of the Large-a Step-Function Limit
Theorem 4 predicts that for stringent SLAs ($a\to\infty$), the imbalance $I(a,h_{0})$ converges to a piecewise constant staircase function, $I_{\infty}(h_{0})$, whose steps are determined solely by the cumulative count of paths, $K(h_{0})$. Figure 2 validates this prediction using a typical large value of $a=10.0$. The validation is robust across a range of large $a$ values; as demonstrated in Appendix G, the simulated curve rapidly converges to the theoretical limit and remains stable, confirming the exponential convergence property. The plot compares the numerically simulated imbalance for a Grid graph at a large but finite strictness ($a=10.0$, solid blue line) with the theoretically derived limit function $I_{\infty}(h_{0})$ (dashed orange line). The agreement is excellent: the simulated curve closely traces the theoretical staircase, with every plateau and every sharp drop occurring at precisely the predicted locations and values. The smooth transitions at the step locations have a narrow width, consistent with the $O(1/a)$ transition width predicted by Proposition 1. This result confirms that our framework can accurately predict the locations of all major performance cliffs for networks operating under strict service requirements.

### v2

@@ I Introduction
We summarize this conceptual comparison in Table I. This paper develops such a framework, centered on an axiomatically-derived QoE-Imbalance metric ($I$) that captures functional fairness through a tunable Sigmoid function ($s(h;a,h_{0})$). Our contributions are threefold:
@@
Ultimately, this work transforms fairness from a static score into a predictable and navigable landscape, offering a new design philosophy for engineering more fundamentally robust and resilient network systems, as Fig. 1 shows.
Figure 1: The conceptual workflow of the proposed fairness sensitivity analysis framework. The framework is structured as a top-to-bottom pipeline, transforming network topology and SLA parameters into a sensitivity landscape. This landscape is then analyzed by a dedicated toolkit, which simultaneously produces three types of outputs: actionable insights, quantitative metrics, and concrete design rules.
@@ II-A An Axiomatic Foundation for Imbalance
Theorem 1 (Uniqueness of the Imbalance Metric).
@@ II-C Analytical Properties and Invariance
A prerequisite for our calculus-based sensitivity analysis is that the objective functions are well-behaved. We formally establish that the metrics $I(a,h_{0})$ and $\bar{s}(a,h_{0})$ are both continuous and differentiable with respect to the SLA parameters $a$ and $h_{0}$ across their entire domain (see Appendix C for formal proofs). This ensures our framework is robust and validates the use of gradients and Hessians.
@@
The proof follows directly from the fact that the exponent in the Sigmoid function, $a(h_{i}-h_{0})$, remains unchanged under both transformations. Since all weights $\{w_{u,v}\}$ are invariant, the resulting metrics $I$ and $\bar{s}$ are also invariant. A full proof is provided in Appendix B. ∎
@@ III-C Second-Order Analysis: The Hessian as a Risk Detector
In the differentiable domain described in Appendix C, the mixed second-order partial derivatives are equal due to Clairaut’s theorem, making the Hessian symmetric. Specifically, the term $|\partial^{2}I/\partial h_{0}^{2}|$ serves as a powerful, quantitative “cliff detector”.
@@ III-D A Unified Methodology for Sensitivity Analysis
    • Area of Robustness (AoR): This metric quantifies the resilience and adaptability of a network to a wide range of service demands. To facilitate a normalized comparison, it is defined as the percentage of the analyzed parameter space that constitutes the Optimal Operating Region $R$:
      $\text{AoR}=\frac{\iint_{R}da\,dh_{0}}{A_{total}}\times 100\%$ (11)
      where $A_{total}$ is the total area of the scanned parameter space. A larger AoR signifies a more robust design.
    • Maximum Curvature Risk (MCR): This metric quantifies the ”sharpness of the cliff” at the edge of acceptable performance. It is defined as the maximum value of the curvature with respect to $h_{0}$ found along the boundary $\partial R$ of the optimal region:
@@
      A lower MCR value indicates a safer design with gentler performance transitions.
@@ IV-B1 Validation of the Small-a Second-Order Law
Theorem 3 predicts that for tolerant SLAs ($a\to 0$), the imbalance $I$ is quadratically proportional to the strictness $a$ ($I\propto a^{2}$), with a coefficient determined by the network’s path-length variance, $\mathrm{Var}(h)$. Fig. 2 provides a direct test of this prediction. As can be seen, the simulation data for all four topologies closely follow a linear trend when $I$ is plotted against $a^{2}$, confirming the quadratic relationship. The slopes of these lines, however, vary dramatically, indicating a strong dependence on the underlying topology. The Path graph, with its extremely dispersed path lengths, is by far the most sensitive, while the ultra-compact Star graph is the least.
@@
*Note: $k_{\text{theory}}=\text{Var}(h)/(8\ln(2)\log_{2}(M))$, with M=N(N1)=2450M=N(N-1)=2450. The factor of $\ln(2)$ in the denominator accounts for using $\log_{2}$ in the definition of $I$ but natural log in the derivation.
@@ IV-B2 Validation of the Large-a Step-Function Limit
Theorem 4 predicts that for stringent SLAs ($a\to\infty$), the imbalance $I(a,h_{0})$ converges to a piecewise constant staircase function, $I_{\infty}(h_{0})$, whose steps are determined solely by the cumulative count of paths, $K(h_{0})$. Fig. 3 validates this prediction using a typical large value of $a=10.0$. The validation is robust across a range of large $a$ values; as demonstrated in Appendix G, the simulated curve rapidly converges to the theoretical limit and remains stable, confirming the exponential convergence property. The plot compares the numerically simulated imbalance for a Grid graph at a large but finite strictness ($a=10.0$, solid blue line) with the theoretically derived limit function $I_{\infty}(h_{0})$ (dashed orange line). The agreement is excellent: the simulated curve closely traces the theoretical staircase, with every plateau and every sharp drop occurring at precisely the predicted locations and values. The smooth transitions at the step locations have a narrow width, consistent with the $O(1/a)$ transition width predicted by Proposition 1. This result confirms that our framework can accurately predict the locations of all major performance cliffs for networks operating under strict service requirements.